OpenStudy (anonymous):

What is the directrix of the parabola with the equation y+3+1/10(x+2) squared?

OpenStudy (jdoe0001):

what the?

OpenStudy (jdoe0001):

so @KelseyCK where's the " = " sign in that parabolic EQUATion?

OpenStudy (jdoe0001):

maybe a quick screenshot would do :)

OpenStudy (anonymous):

There is not an = sign @jdoe0001

OpenStudy (anonymous):

@jdoe0001 I'm sorry the equation is actually y+3=1/10(x+2)squared

OpenStudy (jdoe0001):

\(\large \bf y+3=\cfrac{1}{10}(x+2)^2\quad ?\)

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (jdoe0001):

\(\bf y+3=\cfrac{1}{10}(x+2)^2\implies y+3=\cfrac{(x+2)^2}{10}\implies 10(y+3)=(x+2)^2\) can you see the vertex coordinates of the parabola there?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

I'm just not understanding the question itself and am unsure on how to solve the equation

OpenStudy (dbzfan836):

im sorry, @KelseyCK

OpenStudy (anonymous):

Go away oh my gosh @Dbzfan836